OpenStudy (anonymous):

can someone write these in a expression or equations PLEASE 1) fifteen more than two thirds of a number is thirty 2) two times the sum of five times a number and three 3)twelve minus a number is three more than six squared

OpenStudy (anonymous):

2/3x+15=30 2(5x+3) 12-x+3=6squared(36)

OpenStudy (anonymous):

let x is consider as a number 1) 15+ (2/3)x = 30 2) 2(5x+3) 3) 12- a =3+6^2

OpenStudy (anonymous):

(2/3n) + 15 = 30 2(5n+3) 12 -n = 6^2 +3

OpenStudy (anonymous):

sorry 3 rd one is 3)12- x =3+6^2

OpenStudy (anonymous):

thnk you all!